What is the equilibrium concentration of H2PO4– in a 0.202 M solution of H3PO4(aq)? (Ka1 =...

  1. What is the equilibrium concentration of H2PO4– in a 0.202 M solution of H3PO4(aq)? (Ka1 = 7.5 x 10^-3,Ka2 =6.2x10–8,Ka3 =4.8x10–13)

please list full step by step on how to solve problem.
Answer #1

               H3Po4(aq) ------------------> H2Po4^- (aq)   + H^+ (aq)

I               0.202                                     0                       0

C               -x                                          +x                     +x

E            0.202-x                                     +x                     +x

                          ka1    =   [H2PO4^-][H^+]/[H3PO4]

                         7.5*10^-3    = x*x/(0.202-x)

                         7.5*10^-3*(0.202-x)   = x^2

                               x   = 0.03535

                  [H2PO4^-]   = x    = 0.03535M

the equilibrium concentration of H2PO4–    = 0.03535M